• A firm uses simple exponential smoothing with α = 0.3 to demand forecast. The forecast for the first week of January was 500 units, whereas actual demand turned out to be 450 units.
  • a. Forecast the demand for the second week of January.
  • b. Assume that the actual demand during the second week of January turned out to be 550 units. Forecast the demand up to February third week, assuming the subsequent demands as 475, 450, 470, 525 and 470 units.
1
Expert's answer
2021-02-25T04:10:18-0500

a. Ft = Ft-1 + "\alpha" (At-1 - Ft-1) = 500 + 0.3(450 - 500) = 485 units

b. (475 + 450 + 470 + 525 + 470)÷5 = 478

Ft = 478 + 0.3(550 - 478) = 500 units